### v1

# CO(1–0) imaging reveals 10-kiloparsec molecular gas reservoirs around star-forming galaxies at high redshift

CO(1–0) imaging reveals 10-kiloparsec molecular gas reservoirs around star-forming galaxies at high redshift
Massive, intensely star-forming galaxies at high redshift require a supply of molecular gas from their gas reservoirs, replenished by infall from the surrounding circumgalactic medium, to sustain their immense star-formation rates. However, our knowledge of the extent and morphology of their cold-gas reservoirs is still in its infancy. We present the results of stacking 80 hours of JVLA observations of CO(1–0) emission – which traces the cold molecular gas – in 19 dusty, star-forming galaxies from the AS2VLA survey. The visibility-plane stack reveals extended emission with a half-light radius of kpc, 2–3 more extended than the dust-obscured star formation and more extended than the stellar emission. Similarly, stacking the [C i](1–0) observations for a subsample of our galaxies yields sizes consistent with CO(1–0). The CO(1–0) size is comparable to the [C ii] halos detected around high-redshift star-forming galaxies. The bulk (up to 80%) of molecular gas resides outside the star-forming region; only a small part of their molecular gas reservoir directly contributes to their current star formation. Photon-dissociation region modelling indicates that the extended CO(1–0) emission arises from clumpy, dense clouds rather than smooth, diffuse gas.
Key Words.:
Galaxies: high-redshift — Galaxies: ISM — Submillimeter: galaxies1 Introduction
The majority of star formation beyond redshift took place in massive, dusty star-forming galaxies (DSFGs111Also know as sub-millimetre galaxies – SMGs.) with star-formation rates (SFRs) of a few 100 – 1000 M⊙ yr-1 (e.g., Magnelli et al. 2011; Dudzevičiūtė et al. 2020; Zavala et al. 2021). While some DSFGs are starbursts triggered by major mergers, many appear to be relatively isolated (Gillman et al., 2024; Hodge et al., 2024). The immense SFRs of these isolated galaxies need to be fed by a continuous supply of cold gas from their extended gas reservoirs and the surrounding circumgalactic medium (CGM, e.g., Tumlinson et al. 2017; Faucher-Giguère & Oh 2023) and by minor mergers with gas-rich satellites (e.g., Narayanan et al. 2015).
In local galaxies, cold, molecular gas is primarily traced by the CO(1–0) rotational line, which has an excitation temperature of just 5.5 K and critical density of cm-3. At high redshift, studies of molecular gas have mostly focused on higher-excitation CO rotational emission lines (see, e.g., Hodge & da Cunha 2020; Tacconi et al. 2020; Walter et al. 2020). However, the mid/high-excitation CO emission lines arise from the denser, warmer gas, and are known to be much more compact than the ground-state CO(1–0) transition (e.g., Leroy et al. 2009; Ivison et al. 2011; Emonts et al. 2016), making them unsuitable for tracing cold, diffuse gas.
In contrast to the higher-excitation CO lines, studies of the ground-state CO(1–0) line have been scarce and largely focused on the few brightest galaxies. This is because of the low intrinsic brightness of the CO(1–0) line, a large collecting area and long integrating times are required for achieving the necessary signal-to-noise ratio.
The CO(1–0) emission in a number of high-redshift galaxies has been observed by the Green Bank Telescope (e.g., Hainline et al. 2006; Swinbank et al. 2010; Frayer et al. 2011; Harris et al. 2012; Harrington et al. 2021). However, these spatially unresolved observations do not provide any information about the source size. Early Karl. G. Jansky Very Large Array (JVLA) studies of CO(1–0) emission in bright high-redshift DSFGs (e.g., Ivison et al. 2010, 2011; Riechers et al. 2011b; Thomson et al. 2012) reported half-light radii ranging from a few kpc to 25 kpc (J02399-0136, Thomson et al. 2012), often showing multiple CO(1–0) components. However, some of these results have been challenged by deeper observations: for example, Frias Castillo et al. (2022) showed that the supposedly extended CO(1–0) reservoir with high-velocity wings in the galaxy J13120 (Riechers et al., 2011a) is likely an artefact of low signal-to-noise of the original data. To summarise, the sizes of CO(1–0) reservoirs reported by these early studies might be biased towards the brightest (i.e., most extreme) sources, and suffer from low S/N.
More recently, apparently very extended molecular gas reservoirs have been reported in high-density environments of early protoclusters, such as the Spiderweb (, Emonts et al. 2016) which shows a CO(1–0) reservoirs with a half-light radius of kpc CO(1–0), the -kpc CO(4–3) emission around the star-forming galaxy from Ginolfi et al. (2017), or the claimed super-extended, 200 kpc CO(3–2) emission around a quasar cid_346 (Cicone et al. 2021, but c.f. non-detection in deeper data by Jones et al. 2023).
Further evidence for extended gas around high-redshift DSFGs comes from Atacama Large (Sub)Millimetre Array (ALMA) observations of the [C ii] 158-m line. Extended [C ii] “halos” have been observed around high-redshift DSFGs (e.g., Gullberg et al. 2018; Rybak et al. 2019, 2020) as well as the less massive star-forming galaxies (e.g., Fujimoto et al. 2019, 2020; Ginolfi et al. 2020), sometimes on scales as large as 50 kpc.
The nature of the [C ii] halos is still contested: these could be either low-mass satellite galaxies, outflows from the central galaxy (e.g., Pizzati et al. 2020, 2023) or incoming streams of cold, molecular gas. In particular, interpreting the [C ii] observation is difficult: the [C ii] 158-m emission arises from all gas phases (molecular – H2, atomic – H0, and ionised – H+), rather than just the molecular phase.
The ground-state CO(1–0) emission line remains the most robust probe of diffuse molecular gas. In the recent years, there have been systematic surveys of CO(1–0) in high-redshift DSFGs using JVLA (Frias Castillo et al., 2023; Stanley et al., 2023). These have yielded tentative evidence for extended CO(1–0) emission: Frias Castillo et al. (2023) found that the CO(1–0) extracted from moment-0 maps increases with the aperture size out to radii of 5 (40 kpc). Similarly, using 1-resolution JVLA imaging of 14 galaxies from the z-GAL survey (Cox et al., 2023), Stanley et al. (2023) found CO(1–0) reservoirs of up to 25 kpc across. However, it is unclear how many of their sources are magnified by gravitational lensing, making the interpretation of the measured source sizes difficult.
In this Paper, we report results of stacking 80 hours of CO(1–0) observations from the recently completed AS2VLA222AS2VLA = ALMA, SCUBA-2, VLA. survey (Frias Castillo et al., 2023). Specifically, by using JVLA in the most compact D-array, we combined high surface-brightness sensitivity with ability to recover extended emission. This setup makes them ideally suited for studying potentially very extended, low surface-brightness reservoirs around DSFGs. We identify extended CO(1–0) reservoirs with a scale radius of 4 kpc, indicating that in high-redshift galaxies, the molecular gas extends well beyond the dust and (potentially) stellar emission. The CO(1–0)/[C ii] and dust continuum surface brightness ratios in these reservoirs are consistent with in-situ star-formation, indicating that the extended CO(1–0) and [C ii] halos are a natural extension of the galaxy’s gas reservoirs, rather than residuals of galactic outflows.
Throughout this paper we assume a standard flat CDM cosmology with = 67.8 km s-1 Mpc-1, = 0.310 and = 0.690 (Planck Collaboration et al., 2016).
2 Observations and imaging
2.1 Target sample
Our initial sample consists of 30 sub-mm selected, star-forming galaxies at from the AS2VLA survey (Frias Castillo et al., 2023, Paper I). The sample is selected from ALMA continuum detections of source taken from 4 deg2 of SCUBA-2 850m maps from the S2CLS (Geach et al., 2017) and S2COSMOS (Simpson et al., 2019) surveys. Individual sources were pin-pointed by high-resolution ALMA and SMA imaging (Stach et al., 2018; Hill et al., 2018; Simpson et al., 2020); the 870-m fluxes range between 7.5 and 18.3 mJy. All sources have spectroscopic redshifts based on mid- CO lines and [C i](1–0) (Birkin et al. 2021; Chen et al. 2022a, b, Chapman et al., in prep.). The sources span stellar mass M⊙ and SFR = 400 – 1700 M⊙ yr-1, with a median SFR of 1010 M⊙ yr-1 (Frias Castillo et al., 2023; Liao et al., 2024). Most of them are on the massive end of the main sequence of star-forming galaxies, but several are offset to the starburst regime (SFR/ higher than the main-sequence, Frias Castillo et al. 2023). None of the galaxies is strongly gravitationally lensed; we can thus directly measure their intrinsic sizes.
2.2 JVLA observations
We used JVLA to observe the CO(1–0) emission (rest-frame frequency: = 115.2712 GHz) in thirty DSFGs at . The surveys comprises of the JVLA programs ID:21A-254 (PI: Hodge) and ID:17A-251 (PI: Walter). The observations were taken between March 2021 and December 2023, under good weather conditions in the D-array configuration. The total time per source including overheads was four hours for almost all sources. The exceptions were AS2COS14.1, which was observed for eight hours, and AS2COS44.1 and AS2COS65.1, which were observed for two hours each.
We observed in the K- and Ka-bands, using a contiguous bandwidth of 2 GHz with 2-MHz spectral channels. Depending on the source redshift, the angular resolution ranged from 2.5 to 5.1 (circularised FWHM); the maximum recoverable scale was 44–60. We manually inspected and calibrated the data using the Common Astronomy Software Applications package (Casa, McMullin et al. 2007) versions 6.4.1 and 6.4.3 and the standard JVLA pipeline. The observations of the first 17 galaxies (eleven secure detections) were presented in Frias Castillo et al. (2023). In this work (Paper II), we focus on the 19 sources with secure CO(1–0) detections333For a “secure” detection, we require peak S/N3 in naturally-weighted cleaned maps of CO(1–0) emission. from the completed survey (Paper III, in prep).
2.3 Visibility-plane analysis and stacking
We measure the source sizes directly in the visibility-plane. This approach bypasses several issues with the Cleaned images, including varying beam shape, the fact that cleaned images combine the dirty and clean (Gaussian) beam, and the correlated noise in the image-plane. For illustration, we show the stacked moment-0 image in Fig. 1. Image-plane deconvolution with the imfit task in Casa does not yield a reliable source size.
For the visibility-plane analysis, we process the data as follows: (1) if the wide-band imaging shows a significant (S/N5) continuum, subtract the continuum (constant signal) from the data using the uvcontsub task; (2) select the velocity range within 1 FWHM from the systemic velocity (using linewidths from mid- CO lines, which are detected at a higher S/N); (3) frequency-average the selected data into a single channel; (4) bin the resulting visibilities in radial bins with a size of 2.5, 5, and 10 k. The fields were centred at the ALMA dust continuum peak (accurate to 0.1). We confirmed that the phase-tracking centre is well-aligned with the CO(1–0) emission; we phase-shifted three sources with noticeable offsets (AEG2, CDFN1, CDFN2). For two sources (AS2COS14.1, AS2COS31.1), we found a bright continuum source within the VLA primary beam. We subtracted these companions from the visibilities in Casa, using the ft task.
We fit each binned visibility function with: (1) point-source (constant signal) and (2) an exponential profile. We prefer the exponential profile to the more widely-used Gaussian profile as it provides a more physical description of gas surface brightness distribution in star-forming galaxies 444For the literature source sizes derived using Gaussian profiles, we convert the half-light radius from the Gaussian to exponential profile. (e.g., Leroy et al. 2009; Wang & Lilly 2022) as well as dust and stellar light distribution (e.g., Hodge et al. 2016; Gullberg et al. 2019; Gillman et al. 2024. We utilise the Monte-Carlo Markov chain modelling using the emcee package (Foreman-Mackey et al., 2013). We assumed wide, uniform priors in total flux and half-light radius. For each galaxy, we calculate the Akaike and Bayesian information criteria (AIC, BIC) for each model. The results for individual galaxies are listed in Table 2.
3 Analysis and results
Fits to individual galaxies show that the point-source model (our null hypothesis) is preferred for 15 out of 19 sources. For four sources, exponential profiles with are preferred (see Table 2): these are AS2COS13.1, AS2COS14.1, AS2COS28.1555AS2COS28.1 has a companion offset by 20 kpc, which might contribute to the large apparent source size (Chen et al., 2022a)., and AS2UDS126.0. These four sources do not stand out from the rest of the sample in redshift, SFR or . We hypothesise that our detections of extended emission in these source are due to their lower-than-average redshifts compared to the whole sample (increasing the line signal), which also place the CO(1–0) line at frequencies where JVLA has particularly low noise. In particular, AS2COS13.1 and AS2COS14.1 have the highest S/N out of the entire sample.
We stack the 19 galaxies with CO(1–0) detections and fit them with (1) a point-source, (2) an exponential profile and (3) a combination of an exponential profile + a point-source (to capture a potential combination of a compact and extended component). The exponential model is preferred by both AIC (AIC=27) and BIC (BIC = 27) over the point-source model. The combination of an exponential and a point-source model is not preferred over a single exponential profile.
Changing the bandwidth used for calculation from 2 FWHM to 1 FWHM results in 10% change in the inferred . Similarly, removing baseline longer than 65 k – which might have less robust calibration – results in 10% change. Our results are thus robust with respect to the inclusion of the longest baselines. As our fiducial value, we adopt (3.80.5 kpc at ), derived for a 2FWHM bandwidth using all baselines (Fig. 1, right).
Is the inferred mean source size driven by the four extended sources? As a consistency check, we perform the stacking procedure excluding the four individually extended sources: the preferred model is a single exponential profile with 0.500.10, i.e., almost identical to our fiducial model.
Finally, we stress that the extent of our sources is not due to uncertainties in position and VLA pointing accuracy. The source positions were pin-pointed by sub-arsecond (0.3) ALMA imaging; the VLA positional accuracy is typically 10% of the beam FWHM, i.e. 0.35 for an average observation from our sample. Our tests show that even an extreme 0.5 pointing error yields a stacked source size of 0.09, much smaller than the one inferred from our data.
3.1 Spectral stacking: no signatures for CO(1–0) inflows or outflows
Are the extended CO reservoirs associated with molecular outflows or inflows? Molecular outflows – indicated by high-velocity wings of the line profile - have been invoked to explain extended [C ii] reservoirs around high-redshift galaxies from the ALPINE sample (Ginolfi et al., 2020; Pizzati et al., 2020, 2023); while ALPINE galaxies have SFRs an order-of-magnitude lower than our sample (10–100 M⊙/yr), a similar mechanism might be at play at even more highly star-forming DSFGs.
To search for potential high-velocity components, we stack the spectra of all 19 CO(1–0)-detected galaxies. We adopt Cleaned cubes, made with natural weighting. As the half-light radius of the stacked emission is smaller than the VLA beam, we extract spectra from an aperture equal to the beam size. The stacked spectrum is well-fit by a single Gaussian profile with a line FWHM of 64070 km s-1 (Fig. 2). Although our data is continuum-subtracted, we still find a low positive excess ( 5 Jy, increasing towards higher frequencies) across the entire spectrum. This residual continuum emission is likely due to curvature of the spectrum: at the rest-frame 115 GHz, the combination of free-free and dust thermal emission can lead to a significant spectral curvature. The lack of obvious outflow features indicates that the bulk of CO(1–0) gas is close to the systemic velocity.
3.2 ALMA [CI](1–0) data
We supplement our CO(1–0) data with [C i](1–0) ALMA observations of seven galaxies from our sample which were recently detected in [C i](1–0) line by Frias Castillo et al. (2024). The [C i](1–0) line is considered an alternative tracer of molecular gas, including the “CO-dark” phase (e.g., at low metallicities). The Frias Castillo et al. (2024) ALMA observations had baselines extending up to 140 k, giving a slightly higher angular resolution (2.2-4.0 circularised beam FWHM) than our JVLA data.
Fitting the [C i] visibilities of individual sources, we find that five are consistent with an exponential profile, with half-light radii between 31 and 93 kpc (see Table 2). The two sources that are extended in both CO(1–0) and [C i](1–0) – AS2COS14.1 and AS2UDS126.0 – have consistent sizes in both tracers.
Stacking all the data, we find that the exponential profile with =0.280.10 (2.20.8 kpc) is preferred over the point-source, albeit at lower significance than for CO(1–0) (AIC2.5). The inferred mean [C i](1–0) size is 4025% smaller than the CO(1–0) one, although they are consistent within 1.7. We therefore put an upper limit on the [C i](1–0) size by taking the 95th percentile of the posterior ( 0.42 (3.3 kpc). We thus can not rule out the that [C i](1–0) and CO(1–0) have similar scale lengths.
3.3 PDR modelling
We now infer the thermodynamics of the molecular gas reservoirs using photo-dissociation region (PDR) modelling. Specifically, we estimate the gas density and far-UV irradiation using the PDRToolbox suite of uniform-density PDR models (Kaufman et al., 2006; Pound & Wolfire, 2008). As we will show in Section 4, the CO(1–0)/[C i](1–0) sizes inferred from the uv-plane stacking are consistent with the extended [C ii] and far-IR continuum emission seen in previous studies of DSFGs.
For the far-IR continuum, we assume that the extended component accounts for 13% of the total luminosity (Gullberg et al. 2019, see below). As our sample currently lacks ancillary observations in, e.g., [C ii] 158-m line, we adopt the surface brightness range reported for resolved observations of DSFGs (Gullberg et al., 2018; Rybak et al., 2019; Mitsuhashi et al., 2021): [C ii]/FIR . We assume that 20% of the [C ii] emission arises from ionised gas (as seen in nearby star-forming galaxies, Croxall et al. 2017; Sutter et al. 2019), and apply geometric corrections to the (optically thin) predicted [C ii] and FIR continuum by multiplying them by a factor of two. Finally, we include the [C i](1–0) line, assuming it is co-spatial with CO(1–0).
Figure 4 shows the resulting PDR models, compared to the models for individual galaxies from the AS2VLA sample from Frias Castillo et al. (2024). For the stacked data, we find a median cm-3, (1 erg s-1 cm-2). These are much higher than expected from a supposedly cold, diffuse gas and comparable to values found in Galactic star-forming clouds (e.g., Oberst et al. 2011). The most direct interpretation is that the extended CO(1–0) (and [C ii]) emission arise from in-situ star formation. This interpretation is consistent with results from resolved imaging of H emission – which traces ionised gas – around DSFGs, which revealed extended star formation with a half-light radius of kpc (e.g., Chen et al. 2020; Birkin et al. 2024).
To assess the sensitivity of our conclusions to the various assumptions made in the PDR modelling we have varied these as follows:
What if CO(1–0) is suppressed by CMB? Assuming a 50% suppression (an extreme scenario at , e.g., da Cunha et al. 2013), the inferred far-UV illumination increases by 0.5 dex (blue dotted line).
What if the bulk of [C ii] is arising from ionised gas rather than from the PDRs? Lowering the PDR contribution to the [C ii] flux down to 20% causes the inferred far-UV illumination to increase by almost 1 dex (red dotted line).
What if [C i](1–0) is more compact than CO(1–0) (i.e., the ratio of atomic/molecular carbon increases at large radii)? We assume the [C i](1–0) scale length indicated by the tentative results in Section 3.2 and re-calculate the [C i](1–0) surface brightness at the CO(1–0) half-light radius accordingly. The inferred density increases to cm-3 (green dotted line).
We derive the filling factor – i.e., the number of emitting regions per line-of-sight – by comparing the observed CO(1–0) surface brightness with the PDRToolbox models. Specifically, for each realisation of , , we calculate the corresponding filling factor at the CO(1–0) half-light radius as:
where is the CO(1–0) flux per solid angle predicted by PDRToolbox and is the luminosity distance. Marginalising over and then yields %.
4 Discussion
4.1 Sizes of extended molecular gas reservoirs in high-redshift DSFGs
Figure 5 places our measurement of CO(1–0) sizes in context of other high-redshift observations. Specifically, we show the literature CO(1–0) (Ivison et al., 2011; Thomson et al., 2012; Dannerbauer et al., 2017; Frias Castillo et al., 2022; Stanley et al., 2023) and [C ii] measurements (Gullberg et al., 2018; Rybak et al., 2019; Mitsuhashi et al., 2021). We homogenise the data by adopting the empirical CO excitation ladder from Birkin et al. (2021) and ; for the [C ii] measurements, we assume =30 from Zanella et al. (2018).
Our inferred CO(1–0) size is comparable to CO(1–0) sizes reported in the literature for other highly star-forming galaxy populations, and the [C ii] reservoirs seen around high-redshift DSFGs. The true extent of the molecular gas reservoirs in DSFGs might be even larger than those inferred from our CO(1–0) stacking. First, gas becomes less metal-rich with increasing distance from the centre of the galaxy (e.g., Schimek et al. 2024), causing to increase towards the outskirts. Moreover, the elevated CMB temperature at high redshift can significantly suppress the CO(1–0) emission, particularly at low gas kinetic temperatures and densities (by up to 50%, e.g., da Cunha et al. 2013; Zhang et al. 2016). However, the CMB effect will be more limited if the molecular gas is mostly in the form of higher-density ISM in star-forming regions or satellite galaxies.
4.2 Molecular gas versus dust and stellar emission
How does the size of the extended CO(1–0) emission compare to the far-infrared and near-infrared continuum in DSFGs? In more physical terms, we now compare the molecular gas sizes to star formation (traced by far-IR continuum) and stellar emission (traced by JWST NIRCam imaging). Table 1 lists the sizes of different tracers.
Using high-resolution 870-m ALMA imaging of the SCUBA-2 sources in the UKIDSS/UDS field (from which our sample is partially drawn), Gullberg et al. (2019) found a half-light radius of 1.10.3 kpc, assuming an exponential profile666After applying a correction to account for a faint extended component (Smail et al., 2021).. Somewhat larger sizes (1.80.1 kpc) were inferred for the ALESS sample of DSFGs (Hodge et al., 2016, 2024), which span a comparable redshift range (median ) and 850-m flux ( mJy) to our galaxies. In either case, the 870-m continuum is a factor of a more compact than the CO(1–0) reservoir.
In total, 50–80% of the CO(1–0) emission arises outside the region which contains 90% of the dust continuum emission (adopting the Gullberg et al. 2019 and Hodge et al. 2016 sizes, respectively). The remarkable difference in sizes between the dust continuum and CO(1–0) emission indicates that only a small fraction of the total molecular gas in high-redshift DSFGs contributes directly to their high SFRs. This interpretation is supported by recent studies of HCN/HCO+/HNC emission in DSGFs, which found very low dense-gas fractions (Rybak et al., 2022).
The presence of an extended molecular gas dovetails with the extended low-surface brightness dust continuum reported by Gullberg et al. (2019). After stacking ALMA imaging of 153 DSFGs, Gullberg et al. (2019) reported that 13% of the 870-m continuum arises for an extended component with an assumed radius of 4 kpc, similar to our inferred CO(1–0) size. This extended continuum emission lends further support to the presence of significant metal-enriched gas and reservoirs around high-redshift DSFGs.
As the bulk of molecular gas is far outside the central star-forming region, can the high SFRs of DSFGs be sustained by smooth gas accretion from this reservoirs? We can estimate the required inward streaming velocity by dividing the half-light radius by the median gas depletion time ( Myr, Frias Castillo et al. 2023). We obtain km s-1, comparable to the line-of-sight velocity dispersions seen in high-resolution kinematic studies of DSFGs (10 - 100 km s-1, e.g., Hodge et al. 2012; Rybak et al. 2015; Swinbank et al. 2015; Rizzo et al. 2021, 2023; Birkin et al. 2024. The central star-formation can thus be supported by smooth streaming of gas within the galaxy without major mergers (see, e.g., Narayanan et al. 2015).
For the sizes of stellar emission, we adopt the 4.4-m half-light radius of kpc derived by Gillman et al. (2024) for the AS2COSMOS and AS2UDS DSFGs. A similar value (3.00.3 kpc) was found for the ALESS sample by Hodge et al. (2024). The CO(1–0) emission in our sample is a factor of 1.40.2 and 1.30.2 more extended than 4.4-m sizes from Gillman et al. (2024) and Hodge et al. (2024), respectively. This size discrepancy indicates that in DSFGs, the cold molecular gas appears to extend beyond the stellar emission of the galaxy.
Tracer | [kpc] | Reference |
CO(1–0) | 3.80.5 | This work |
[C i](1–0) | 3.3 | This work |
870 m | 1.10.3 | Gullberg et al. (2019) |
870 m | 1.80.1 | Hodge et al. (2016) |
4.4 m | 2.70.2 | Gillman et al. (2024) |
4.4 m | 3.00.3 | Hodge et al. (2024) |
The discrepancy in the spatial extent of molecular gas versus star formation and stars might be even more pronounced than indicated from the CO(1–0) and 4.4-m imaging. First, molecular gas might be more extended than CO(1–0) emission, due to higher and increased CMB attenuation in the low-density outskirts of galaxies. Second, if the dust temperature decreases towards larger radii (e.g., Calistro Rivera et al. 2018) star-formation will be more centrally concentrated than the observed-frame 870-m continuum. Finally, stellar component is likely more compact than the near-infrared continuum; high obscuration in the central region will “inflate” the apparent near-infrared sizes compared to the actual stellar component (Sorba & Sawicki, 2018; Popping et al., 2022; Smail et al., 2023; Gillman et al., 2024).
How do the relative sizes of molecular gas, stars and dust continuum in high-redshift DSFGs compare to other galaxy populations? At , spiral galaxies have molecular gas reservoirs with half-light radii of 1–5 kpc (e.g., Regan et al. 2001; Leroy et al. 2009). The CO(1–0)/(2–1) emission in spiral galaxies is typically smaller than the stellar emission (e.g., Regan et al. 2001 find a mean ratio of 0.880.14). In contrast, the ultraluminous infrared galaxies (ULIRGs) have very compact molecular gas reservoirs (typically less 1 kpc, e.g., Bellocchi et al. 2022), a factor of a few smaller than their stellar components. The larger extent of molecular gas compared to the stellar component in DSFGs suggests an inside-out growth (as implied by the small dust continuum sizes) or a presence of satellite galaxies.
4.3 Comparison to molecular gas in hydrodynamical simulations
How do our findings compare to predictions from simulations? Diffuse molecular gas has been notoriously difficult to produce in large-volume hydrodynamical simulations, as convergence requires very high (pc-scale) resolution (e.g., Faucher-Giguère & Oh 2023).
For a comparison in Fig. 5, we take data from the SIMBA cosmological simulation (Davé et al., 2019). SIMBA successfully reproduces key parameters of high-redshift DSFGs (Lovell et al., 2021). We focus the snapshot from the default 100-cMpc box. We select galaxies with M⊙, i.e., comparable in mass to our sample. Specifically, we consider the molecular gas masses and the (total) gas half-light radii from the SIMBA catalogue. We note that SIMBA does not predict H2 directly; rather, the molecular gas is predicted by post-processing the simulation data (Davé et al., 2019). The gas sizes are determined by averaging two-dimensional projections along the three axes.
As shown in Fig. 5, SIMBA galaxies with the highest molecular-gas masses have sizes and mean surface densities comparable to our sample. However, unlike the observed galaxies, SIMBA galaxies tend to become more compact with increasing SFR. This discrepancy is likely a result of the sub-grid star-formation prescription used in SIMBA, which implements a Kennicutt-Schmidt law (i.e., star-formation rate scales with the gas surface density, Davé et al. 2019).
4.4 The nature and thermodynamics of the extended CO(1–0) reservoirs
What is the nature of the extended CO(1–0) emission around DSFGs? Are they smooth extensions of the galaxy’s gas reservoirs, tidal features induced by galaxy mergers, or individual faint satellites? Discriminating between different scenarios requires deep, sub-arcsecond observations of individual galaxies, while preserving sensitivity to large spatial scales. Unfortunately, our stacking averages the data in the azimuthal direction - i.e., smearing emission from compact satellites.
However, we can get significant insights using results from the PDR modelling. As shown in Section 3.3, the inferred gas density and far-UV illumination are comparable to those in Galactic star-forming clouds. Conversely, the filling factor (%) is smaller than the disks of nearby galaxies, in which it ranges between few per cent to factor unity (e.g., Wolfire et al. 1990; Walsh et al. 2002; Luhman et al. 2003).
The combination of high , and low indicates that molecular gas resides in compact, dense clouds, that are distributed sparsely on the outskirts of galaxies. Alternatively, the clouds might be concentrated in satellites with small angular separation from the host (although we note the lack of companions at 1 separation in JWST NIRCam imaging of DSFGs in Gillman et al. 2024; Hodge et al. 2024). In either scenario, the molecular gas on the galaxy outskirts is significantly clumpy rather than distributed in a smooth, diffuse “halo”.
5 Conclusions
We have presented results of stacking CO(1–0) observations of a sample of 19 high-redshift DSFGs. Comprising 80 hours in the most compact JVLA array configuration, this is the deepest such study to date. Our main findings are:
The stacked -plane data show an extended exponential profile with a half-light radius of 3.80.5 kpc. The inferred size is comparable to extended CO(1–0) and [C ii] reservoirs previously reported around high-redshift DSFGs.
Four galaxies have clearly extended CO(1–0) emission, with half-light radii between 3.4–7.2 kpc.
Applying the same procedure to ALMA [C i](1–0) observations, we find that [C i](1–0) appears marginally more compact (2.20.8 kpc) than CO(1–0), although both sizes are consistent within 2.
Using PDR modelling to infer physical conditions in the extended CO(1–0) reservoirs, we find a density of cm-3 and a far-UV field of , and a filling factor of 1%. These results indicate that the extended CO (and [C i] and [C ii]) emission arises from sparsely distributed star-forming clouds.
The inferred sizes of CO(1–0) reservoir are significantly larger than the 870-m dust continuum, indicating that the bulk of molecular gas (50–80%) resides outside the FIR-bright star-forming region. CO(1–0) is somewhat larger than the 4.4-m stellar emission.
Ultimately, understanding the nature of extended molecular gas reservoirs around high-redshift galaxies will require a new class of large, single-dish telescopes, such as the proposed 50-metre Atacama Large Aperture Submillimeter Telescope (AtLAST) (Mroczkowski et al., 2024; Lee et al., 2024).
Acknowledgements.
M.R. is supported by the NWO Veni project ”Under the lens” (VI.Veni.202.225). J.A.H. acknowledges support from the ERC Consolidator Grant 101088676 (VOYAJ). I.R.S. and A.M.S. acknowledge STFC grant ST/X001075/1. E.F.-J.A. acknowledges support from UNAM-PAPIIT project IA102023, and from CONAHCyT Ciencia de Frontera project ID: CF-2023-I-506. C.-C.C. acknowledges support from the National Science and Technology Council of Taiwan (111-2112M-001-045-MY3), as well as Academia Sinica through the Career Development Award (AS-CDA-112-M02). The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc. This paper makes use of the following ALMA data: ADS/JAO.ALMA #2021.1.01324.S. ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC (Canada), MOST and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ. The authors are thankful for the assistance from Allegro, the European ALMA Regional Center node in the Netherlands. We thank D. Narayanan, E. Pizzati and A. Schimek for sharing insights from theoretical models and simulations.References
- Bellocchi et al. (2022) Bellocchi, E., Pereira-Santaella, M., Colina, L., et al. 2022, A&A, 664, A60
- Birkin et al. (2024) Birkin, J. E., Puglisi, A., Swinbank, A. M., et al. 2024, MNRAS, 531, 61
- Birkin et al. (2021) Birkin, J. E., Weiss, A., Wardlow, J. L., et al. 2021, MNRAS, 501, 3926
- Calistro Rivera et al. (2018) Calistro Rivera, G., Hodge, J. A., Smail, I., et al. 2018, ApJ, 863, 56
- Chen et al. (2020) Chen, C.-C., Harrison, C. M., Smail, I., et al. 2020, A&A, 635, A119
- Chen et al. (2022a) Chen, C.-C., Liao, C.-L., Smail, I., et al. 2022a, ApJ, 929, 159
- Chen et al. (2022b) Chen, C.-C., Liao, C.-L., Smail, I., et al. 2022b, ApJ, 929, 159
- Cicone et al. (2021) Cicone, C., Mainieri, V., Circosta, C., et al. 2021, A&A, 654, L8
- Cox et al. (2023) Cox, P., Neri, R., Berta, S., et al. 2023, A&A, 678, A26
- Croxall et al. (2017) Croxall, K. V., Smith, J. D., Pellegrini, E., et al. 2017, ApJ, 845, 96
- da Cunha et al. (2013) da Cunha, E., Groves, B., Walter, F., et al. 2013, ApJ, 766, 13
- Dannerbauer et al. (2017) Dannerbauer, H., Lehnert, M. D., Emonts, B., et al. 2017, A&A, 608, A48
- Davé et al. (2019) Davé, R., Anglés-Alcázar, D., Narayanan, D., et al. 2019, MNRAS, 486, 2827
- Dudzevičiūtė et al. (2020) Dudzevičiūtė, U., Smail, I., Swinbank, A. M., et al. 2020, Monthly Notices of the Royal Astronomical Society, 494, 3828–3860
- Emonts et al. (2016) Emonts, B. H. C., Lehnert, M. D., Villar-Martín, M., et al. 2016, Science, 354, 1128
- Faucher-Giguère & Oh (2023) Faucher-Giguère, C.-A. & Oh, S. P. 2023, ARA&A, 61, 131
- Foreman-Mackey et al. (2013) Foreman-Mackey, D., Hogg, D. W., Lang, D., & Goodman, J. 2013, PASP, 125, 306
- Frayer et al. (2011) Frayer, D. T., Harris, A. I., Baker, A. J., et al. 2011, ApJ, 726, L22
- Frias Castillo et al. (2023) Frias Castillo, M., Hodge, J., Rybak, M., et al. 2023, ApJ, 945, 128
- Frias Castillo et al. (2022) Frias Castillo, M., Rybak, M., Hodge, J., et al. 2022, ApJ, 930, 35
- Frias Castillo et al. (2024) Frias Castillo, M., Rybak, M., Hodge, J. A., et al. 2024, arXiv e-prints, arXiv:2404.05596
- Fujimoto et al. (2019) Fujimoto, S., Ouchi, M., Ferrara, A., et al. 2019, ApJ, 887, 107
- Fujimoto et al. (2020) Fujimoto, S., Silverman, J. D., Bethermin, M., et al. 2020, ApJ, 900, 1
- Geach et al. (2017) Geach, J. E., Dunlop, J. S., Halpern, M., et al. 2017, MNRAS, 465, 1789
- Gillman et al. (2024) Gillman, S., Smail, I., Gullberg, B., et al. 2024, arXiv e-prints, arXiv:2406.03544
- Ginolfi et al. (2020) Ginolfi, M., Jones, G. C., Béthermin, M., et al. 2020, A&A, 633, A90
- Ginolfi et al. (2017) Ginolfi, M., Maiolino, R., Nagao, T., et al. 2017, MNRAS, 468, 3468
- Gullberg et al. (2019) Gullberg, B., Smail, I., Swinbank, A. M., et al. 2019, MNRAS, 490, 4956
- Gullberg et al. (2018) Gullberg, B., Swinbank, A. M., Smail, I., et al. 2018, ApJ, 859, 12
- Hainline et al. (2006) Hainline, L. J., Blain, A. W., Greve, T. R., et al. 2006, ApJ, 650, 614
- Harrington et al. (2021) Harrington, K. C., Weiss, A., Yun, M. S., et al. 2021, ApJ, 908, 95
- Harris et al. (2012) Harris, A. I., Baker, A. J., Frayer, D. T., et al. 2012, ApJ, 752, 152
- Hill et al. (2018) Hill, R., Chapman, S. C., Scott, D., et al. 2018, MNRAS, 477, 2042
- Hodge et al. (2012) Hodge, J. A., Carilli, C. L., Walter, F., et al. 2012, ApJ, 760, 11
- Hodge & da Cunha (2020) Hodge, J. A. & da Cunha, E. 2020, Royal Society Open Science, 7, 200556
- Hodge et al. (2024) Hodge, J. A., da Cunha, E., Kendrew, S., et al. 2024, arXiv e-prints, arXiv:2407.15846
- Hodge et al. (2016) Hodge, J. A., Swinbank, A. M., Simpson, J. M., et al. 2016, ApJ, 833, 103
- Ivison et al. (2011) Ivison, R. J., Papadopoulos, P. P., Smail, I., et al. 2011, MNRAS, 412, 1913
- Ivison et al. (2010) Ivison, R. J., Smail, I., Papadopoulos, P. P., et al. 2010, MNRAS, 404, 198
- Jones et al. (2023) Jones, G. C., Maiolino, R., Carniani, S., et al. 2023, MNRAS, 522, 275
- Kaufman et al. (2006) Kaufman, M. J., Wolfire, M. G., & Hollenbach, D. J. 2006, ApJ, 644, 283
- Lee et al. (2024) Lee, M. M., Schimek, A., Cicone, C., et al. 2024, arXiv e-prints, arXiv:2403.00924
- Leroy et al. (2009) Leroy, A. K., Walter, F., Bigiel, F., et al. 2009, AJ, 137, 4670
- Liao et al. (2024) Liao, C.-L., Chen, C.-C., Wang, W.-H., et al. 2024, ApJ, 961, 226
- Lovell et al. (2021) Lovell, C. C., Geach, J. E., Davé, R., Narayanan, D., & Li, Q. 2021, MNRAS, 502, 772
- Luhman et al. (2003) Luhman, M. L., Satyapal, S., Fischer, J., et al. 2003, ApJ, 594, 758
- Magnelli et al. (2011) Magnelli, B., Elbaz, D., Chary, R. R., et al. 2011, A&A, 528, A35
- McMullin et al. (2007) McMullin, J. P., Waters, B., Schiebel, D., Young, W., & Golap, K. 2007, in Astronomical Society of the Pacific Conference Series, Vol. 376, Astronomical Data Analysis Software and Systems XVI, ed. R. A. Shaw, F. Hill, & D. J. Bell, 127
- Mitsuhashi et al. (2021) Mitsuhashi, I., Matsuda, Y., Smail, I., et al. 2021, ApJ, 907, 122
- Mroczkowski et al. (2024) Mroczkowski, T., Gallardo, P. A., Timpe, M., et al. 2024, arXiv e-prints, arXiv:2402.18645
- Narayanan et al. (2015) Narayanan, D., Turk, M., Feldmann, R., et al. 2015, Nature, 525, 496
- Oberst et al. (2011) Oberst, T. E., Parshley, S. C., Nikola, T., et al. 2011, ApJ, 739, 100
- Pizzati et al. (2020) Pizzati, E., Ferrara, A., Pallottini, A., et al. 2020, MNRAS, 495, 160
- Pizzati et al. (2023) Pizzati, E., Ferrara, A., Pallottini, A., et al. 2023, MNRAS, 519, 4608
- Planck Collaboration et al. (2016) Planck Collaboration, Ade, P. A. R., Aghanim, N., et al. 2016, A&A, 594, A13
- Popping et al. (2022) Popping, G., Pillepich, A., Calistro Rivera, G., et al. 2022, MNRAS, 510, 3321
- Pound & Wolfire (2008) Pound, M. W. & Wolfire, M. G. 2008, in Astronomical Society of the Pacific Conference Series, Vol. 394, Astronomical Data Analysis Software and Systems XVII, ed. R. W. Argyle, P. S. Bunclark, & J. R. Lewis, 654
- Regan et al. (2001) Regan, M. W., Thornley, M. D., Helfer, T. T., et al. 2001, ApJ, 561, 218
- Riechers et al. (2011a) Riechers, D. A., Carilli, L. C., Walter, F., et al. 2011a, ApJ, 733, L11
- Riechers et al. (2011b) Riechers, D. A., Hodge, J., Walter, F., Carilli, C. L., & Bertoldi, F. 2011b, ApJ, 739, L31
- Rizzo et al. (2023) Rizzo, F., Roman-Oliveira, F., Fraternali, F., et al. 2023, A&A, 679, A129
- Rizzo et al. (2021) Rizzo, F., Vegetti, S., Fraternali, F., Stacey, H. R., & Powell, D. 2021, MNRAS, 507, 3952
- Rybak et al. (2019) Rybak, M., Calistro Rivera, G., Hodge, J. A., et al. 2019, ApJ, 876, 112
- Rybak et al. (2022) Rybak, M., Hodge, J. A., Greve, T. R., et al. 2022, A&A, 667, A70
- Rybak et al. (2020) Rybak, M., Hodge, J. A., Vegetti, S., et al. 2020, MNRAS, 494, 5542
- Rybak et al. (2015) Rybak, M., Vegetti, S., McKean, J. P., Andreani, P., & White, S. D. M. 2015, MNRAS, 453, L26
- Schimek et al. (2024) Schimek, A., Decataldo, D., Shen, S., et al. 2024, A&A, 682, A98
- Simpson et al. (2020) Simpson, J. M., Smail, I., Dudzevičiūtė, U., et al. 2020, MNRAS, 495, 3409
- Simpson et al. (2019) Simpson, J. M., Smail, I., Swinbank, A. M., et al. 2019, ApJ, 880, 43
- Smail et al. (2023) Smail, I., Dudzevičiūtė, U., Gurwell, M., et al. 2023, ApJ, 958, 36
- Smail et al. (2021) Smail, I., Dudzevičiūtė, U., Stach, S. M., et al. 2021, MNRAS, 502, 3426
- Sorba & Sawicki (2018) Sorba, R. & Sawicki, M. 2018, MNRAS, 476, 1532
- Stach et al. (2018) Stach, S. M., Smail, I., Swinbank, A. M., et al. 2018, ApJ, 860, 161
- Stanley et al. (2023) Stanley, F., Jones, B. M., Riechers, D. A., et al. 2023, ApJ, 945, 24
- Sutter et al. (2019) Sutter, J., Dale, D. A., Croxall, K. V., et al. 2019, ApJ, 886, 60
- Swinbank et al. (2015) Swinbank, A. M., Dye, S., Nightingale, J. W., et al. 2015, ApJ, 806, L17
- Swinbank et al. (2010) Swinbank, A. M., Smail, I., Chapman, S. C., et al. 2010, MNRAS, 405, 234
- Tacconi et al. (2020) Tacconi, L. J., Genzel, R., & Sternberg, A. 2020, ARA&A, 58, 157
- Thomson et al. (2012) Thomson, A. P., Ivison, R. J., Smail, I., et al. 2012, MNRAS, 425, 2203
- Tumlinson et al. (2017) Tumlinson, J., Peeples, M. S., & Werk, J. K. 2017, ARA&A, 55, 389
- Walsh et al. (2002) Walsh, W., Beck, R., Thuma, G., et al. 2002, A&A, 388, 7
- Walter et al. (2020) Walter, F., Carilli, C., Neeleman, M., et al. 2020, ApJ, 902, 111
- Wang & Lilly (2022) Wang, E. & Lilly, S. J. 2022, ApJ, 927, 217
- Wolfire et al. (1990) Wolfire, M. G., Tielens, A. G. G. M., & Hollenbach, D. 1990, ApJ, 358, 116
- Zanella et al. (2018) Zanella, A., Daddi, E., Magdis, G., et al. 2018, MNRAS, 481, 1976
- Zavala et al. (2021) Zavala, J. A., Casey, C. M., Manning, S. M., et al. 2021, ApJ, 909, 165
- Zhang et al. (2016) Zhang, Z.-Y., Papadopoulos, P. P., Ivison, R. J., et al. 2016, Royal Society Open Science, 3, 160025
Appendix A Inferred source sizes
Table 2 lists the source sizes inferred from the MCMC modelling, using 5-k radial bin size. For galaxies for which the extended source is not preferred by both AIC and BIC, we give upper limit on the source sized, based on the 95th percentile from the posterior of the exponential model.
Source ID | CO(1–0) | [C i](1–0) | |||
[arcsec] | [kpc] | [arcsec] | [kpc] | ||
AS2COS1.1 | 4.625 | 2.62 | 17 | 1.36 | 9 |
AS2COS8.1 | 3.581 | 1.05 | 8 | 0.75 | 6 |
AS2COS13.1 | 2.608 | 0.41 | 2.9 | — | — |
AS2COS14.1 | 2.921 | 0.74 | 5.8 | 0.98 | 7.7 |
AS2COS23.1 | 4.341 | 1.89 | 13 | 1.110.44 | 7.63.0 |
AS2COS28.1 | 3.097 | 0.66 | 5.1 | — | — |
AS2COS31.1 | 3.643 | 5.26 | 39 | — | — |
AS2COS54.1 | 3.174 | 1.60 | 12 | — | — |
AS2COS65.1 | 2.414 | 0.68 | 5.6 | 0.380.11 | 3.10.9 |
AS2COS66.1 | 3.247 | 2.09 | 17 | — | — |
AS2COS139.1 | 2.219 | 1.08 | 8 | — | — |
AS2UDS10.0 | 3.169 | 3.65 | 28 | — | — |
AS2UDS11.0 | 4.073 | 2.37 | 38 | — | — |
AS2UDS12.0 | 2.520 | 0.80 | 6.6 | 0.400.19 | 3.31.6 |
AS2UDS126.0 | 2.436 | 0.97 | 8.0 | 0.640.28 | 5.32.3 |
AEG2 | 3.668 | 1.43 | 12 | — | — |
CDFN1 | 3.149 | 0.80 | 6 | — | — |
CDFN2 | 4.422 | 1.35 | 9 | — | — |

### v2

@@
 CO(1–0) imaging reveals 10-kiloparsec molecular gas reservoirs around star-forming galaxies at high redshift
+11email: mrybak@strw.leidenuniv.nl 22institutetext: Faculty of Electrical Engineering, Mathematics and Computer Science, Delft University of Technology, Mekelweg 4, 2628 CD Delft, The Netherlands 33institutetext: SRON – Netherlands Institute for Space Research, Niels Bohrweg 4, 2333 CA Leiden, The Netherlands 44institutetext: Centre for Extragalactic Astronomy, Department of Physics, Durham University, South Road, Durham DH1 3LE, UK 55institutetext: German Aerospace Center (DLR), Institute of Communications and Navigation, Wessling, Germany 66institutetext: European Southern Observatory (ESO), Karl-Schwarzschild-Straße 2, 85748 Garching bei München, Germany 77institutetext: Department of Physics and Atmospheric Science, Dalhousie University, Halifax, NS, B3H 4R2, Canada 88institutetext: Academia Sinica Institute of Astronomy and Astrophysics (ASIAA), No.
+1, Section 4, Roosevelt Road, Taipei 106216, Taiwan 99institutetext: International Centre for Radio Astronomy Research, University of Western Australia, 7 Fairway, Crawley, Perth, Australia 1010institutetext: ARC Centre of Excellence for All Sky Astrophysics in 3 Dimensions (ASTRO 3D), Australia 1111institutetext: Cosmic Dawn Center (DAWN), Denmark 1212institutetext: Instituto de Astrofísica de Canarias, Vía Láctea, 39020 La Laguna (Tenerife), Spain 1313institutetext: Instituto de Radioastronomía y Astrofísica, Universidad Nacional Autónoma de México, Antigua Carretera a Pátzcuaro # 8701, Ex-Hda.
+San José de la Huerta, Morelia, Michoacán, México C.P.
+58089 1414institutetext: National Radio Astronomy Observatory, Charlottesville, VA, USA 1515institutetext: Department of Physics and Astronomy, University of British Columbia, 6224 Agricultural Road, Vancouver, BC V6T 1Z1, Canada 1616institutetext: Max–Planck Institut für Astronomie, Königstuhl 17, 69117 Heidelberg, Germany
 CO(1–0) imaging reveals 10-kiloparsec molecular gas reservoirs around star-forming galaxies at high redshift
@@
 However, our knowledge of the extent and morphology of their cold-gas reservoirs is still in its infancy.
-We present the results of stacking 80 hours of JVLA observations of CO(1–0) emission – which traces the cold molecular gas – in 19 dusty, star-forming galaxies from the AS2VLA survey.
-The visibility-plane stack reveals extended emission with a half-light radius of kpc, 2–3 more extended than the dust-obscured star formation and more extended than the stellar emission.
-Similarly, stacking the [C i](1–0) observations for a subsample of our galaxies yields sizes consistent with CO(1–0).
-The CO(1–0) size is comparable to the [C ii] halos detected around high-redshift star-forming galaxies.
-The bulk (up to 80%) of molecular gas resides outside the star-forming region; only a small part of their molecular gas reservoir directly contributes to their current star formation.
-Photon-dissociation region modelling indicates that the extended CO(1–0) emission arises from clumpy, dense clouds rather than smooth, diffuse gas.
+We present the results of stacking 80 hours of JVLA observations of CO(1–0) emission – which traces the cold molecular gas – in nineteen dusty, star-forming galaxies from the AS2VLA survey.
+The visibility-plane stack reveals extended emission with a half-light radius of kpc, 2–3 more extended than the dust-obscured star formation and more extended than the stellar emission revealed by JWST.
+Stacking the [C i](1–0) observations for ten galaxies from our parent sample yields a half-light radius 2.6 kpc, marginally smaller than CO(1–0).
+The CO(1–0) size is also comparable to the [C ii] halos detected around high-redshift star-forming galaxies, suggesting these arise from molecular gas.
+Photo-dissociation region modelling indicates that the extended CO(1–0) emission arises from clumpy, dense clouds rather than smooth, diffuse gas.
+Our results show that the bulk (up to 80%) of molecular gas resides outside the star-forming region; with only a small part directly contributing to their current star formation.
 Key Words.:
 Galaxies: high-redshift — Galaxies: ISM — Submillimeter: galaxies1 Introduction
-The majority of star formation beyond redshift took place in massive, dusty star-forming galaxies (DSFGs111Also know as sub-millimetre galaxies – SMGs.) with star-formation rates (SFRs) of a few 100 – 1000 M⊙ yr-1 (e.g., Magnelli et al.
+The majority of star formation beyond redshift took place in massive, dusty star-forming galaxies (DSFGs111Also know as sub-millimetre galaxies (SMGs).) with star-formation rates (SFRs) of a few 100 – 1000 M⊙ yr-1 (e.g., Magnelli et al.
 2011; Dudzevičiūtė et al.
@@
 2021).
-While some DSFGs are starbursts triggered by major mergers, many appear to be relatively isolated (Gillman et al., 2024; Hodge et al., 2024).
+While some DSFGs are starbursts triggered by major mergers, many appear to be relatively isolated (Gillman et al., 2024; Hodge et al., 2025).
 The immense SFRs of these isolated galaxies need to be fed by a continuous supply of cold gas from their extended gas reservoirs and the surrounding circumgalactic medium (CGM, e.g., Tumlinson et al.
-2017; Faucher-Giguère & Oh 2023) and by minor mergers with gas-rich satellites (e.g., Narayanan et al.
+2017; Faucher-Giguère & Oh 2023) or by minor mergers with gas-rich satellites (e.g., Narayanan et al.
 2015).
@@
 2020; Walter et al.
-2020).
-However, the mid/high-excitation CO emission lines arise from the denser, warmer gas, and are known to be much more compact than the ground-state CO(1–0) transition (e.g., Leroy et al.
+2020; Birkin et al.
+2021).
+However, the mid/high-excitation CO emission lines arise from the denser, warmer gas, and are expected to be much more compact than the ground-state CO(1–0) transition (e.g., Leroy et al.
 2009; Ivison et al.
@@
 2016), making them unsuitable for tracing cold, diffuse gas.
-In contrast to the higher-excitation CO lines, studies of the ground-state CO(1–0) line have been scarce and largely focused on the few brightest galaxies.
-This is because of the low intrinsic brightness of the CO(1–0) line, a large collecting area and long integrating times are required for achieving the necessary signal-to-noise ratio.
+In contrast to the higher-excitation CO lines, studies of the ground-state CO(1–0) line in high-redshift galaxies have been scarce and largely focused on the few brightest sources.
+This is because of the low intrinsic brightness of the CO(1–0) line, which requires a large collecting area and long integrating times to achieve the necessary signal-to-noise ratio (S/N).
 The CO(1–0) emission in a number of high-redshift galaxies has been observed by the Green Bank Telescope (e.g., Hainline et al.
@@
 2010; Frayer et al.
-2011; Harris et al.
+2011, 2018; Harris et al.
 2012; Harrington et al.
@@
 2012) reported half-light radii ranging from a few kpc to 25 kpc (J02399-0136, Thomson et al.
-2012), often showing multiple CO(1–0) components.
+2012; Frayer et al.
+2018), often showing multiple CO(1–0) components.
 However, some of these results have been challenged by deeper observations: for example, Frias Castillo et al.
-(2022) showed that the supposedly extended CO(1–0) reservoir with high-velocity wings in the galaxy J13120 (Riechers et al., 2011a) is likely an artefact of low signal-to-noise of the original data.
+(2022) showed that the supposedly extended CO(1–0) reservoir with high-velocity wings in the galaxy J13120+4242 (Riechers et al., 2011a) is likely an artefact of low signal-to-noise of the original data.
 To summarise, the sizes of CO(1–0) reservoirs reported by these early studies might be biased towards the brightest (i.e., most extreme) sources, and suffer from low S/N.
@@
 2016) which shows a CO(1–0) reservoirs with a half-light radius of kpc CO(1–0), the -kpc CO(4–3) emission around the star-forming galaxy from Ginolfi et al.
-(2017), or the claimed super-extended, 200 kpc CO(3–2) emission around a quasar cid_346 (Cicone et al.
-2021, but c.f.
-non-detection in deeper data by Jones et al.
-2023).
+(2017), or the claimed super-extended, 200 kpc CO(3–2) emission around a quasar cid_346 (Cicone et al., 2021), which was however not confirmed with deeper data (Jones et al., 2023)).
 Further evidence for extended gas around high-redshift DSFGs comes from Atacama Large (Sub)Millimetre Array (ALMA) observations of the [C ii] 158-m line.
@@
 2018; Rybak et al.
-2019, 2020) as well as the less massive star-forming galaxies (e.g., Fujimoto et al.
+2019, 2020) as well as around less massive star-forming galaxies (e.g., Fujimoto et al.
 2019, 2020; Ginolfi et al.
-2020), sometimes on scales as large as 50 kpc.
+2020; Ikeda et al.
+2025), sometimes on scales as large as 50 kpc.
 The nature of the [C ii] halos is still contested: these could be either low-mass satellite galaxies, outflows from the central galaxy (e.g., Pizzati et al.
@@
 (2023) found CO(1–0) reservoirs of up to 25 kpc across.
-However, it is unclear how many of their sources are magnified by gravitational lensing, making the interpretation of the measured source sizes difficult.
+However, as about 50% of the Stanley et al.
+(2023) sources are gravitationally lensed, the interpretation of the measured source sizes is not straightforward.
 In this Paper, we report results of stacking 80 hours of CO(1–0) observations from the recently completed AS2VLA222AS2VLA = ALMA, SCUBA-2, VLA.
 survey (Frias Castillo et al., 2023).
-Specifically, by using JVLA in the most compact D-array, we combined high surface-brightness sensitivity with ability to recover extended emission.
-This setup makes them ideally suited for studying potentially very extended, low surface-brightness reservoirs around DSFGs.
-We identify extended CO(1–0) reservoirs with a scale radius of 4 kpc, indicating that in high-redshift galaxies, the molecular gas extends well beyond the dust and (potentially) stellar emission.
-The CO(1–0)/[C ii] and dust continuum surface brightness ratios in these reservoirs are consistent with in-situ star-formation, indicating that the extended CO(1–0) and [C ii] halos are a natural extension of the galaxy’s gas reservoirs, rather than residuals of galactic outflows.
+Specifically, using the JVLA in the most compact D-array and aiming at high surface-brightness sensitivity, provides the ideal set-up to study potentially very extended, low surface-brightness cold gas reservoirs around high-redshift DSFGs.
+This Paper is structured as follows: in Section 2, we describe our sample selection, details of observations, and the visibility-plane analysis and stacking procedure.
+We then look for potential cold-gas outflows (Section 3.1), present the ancillary [C i](1–0) data, and use PDR modelling to infer the thermodynamics of extended gas reservoirs (3.3).
+Finally, we compare our inferred CO(1–0) sizes to observations of dust, [C ii], and stellar emission (Section 4.1 and 4.2) and predictions from hydrodynamical simulations (4.3).
 Throughout this paper we assume a standard flat CDM cosmology with = 67.8 km s-1 Mpc-1, = 0.310 and = 0.690 (Planck Collaboration et al., 2016).
@@
 2.1 Target sample
-Our initial sample consists of 30 sub-mm selected, star-forming galaxies at from the AS2VLA survey (Frias Castillo et al., 2023, Paper I).
+Our initial sample consists of 30 sub-millimetre selected, star-forming galaxies at from the AS2VLA survey (Frias Castillo et al., 2023, Paper I).
 The sample is selected from ALMA continuum detections of source taken from 4 deg2 of SCUBA-2 850m maps from the S2CLS (Geach et al., 2017) and S2COSMOS (Simpson et al., 2019) surveys.
@@
 The sources span stellar mass M⊙ and SFR = 400 – 1700 M⊙ yr-1, with a median SFR of 1010 M⊙ yr-1 (Frias Castillo et al., 2023; Liao et al., 2024).
-Most of them are on the massive end of the main sequence of star-forming galaxies, but several are offset to the starburst regime (SFR/ higher than the main-sequence, Frias Castillo et al.
+and SFR were derived by Dudzevičiūtė et al.
+(2020); Simpson et al.
+(2020) using the high-redshift version of the MagPhys spectral energy distribution modelling code (da Cunha et al., 2015; Battisti et al., 2019).
+Most of them are on the massive end of the main sequence of star-forming galaxies, but several are offset to the starburst regime (SFR/ than the main-sequence, Frias Castillo et al.
 2023).
-None of the galaxies is strongly gravitationally lensed; we can thus directly measure their intrinsic sizes.
+None of the galaxies is strongly gravitationally lensed, as confirmed by ALMA 0.3”-resolution continuum imaging (Stach et al., 2018; Simpson et al., 2020).
+We can thus directly measure their intrinsic sizes.
 2.2 JVLA observations
-We used JVLA to observe the CO(1–0) emission (rest-frame frequency: = 115.2712 GHz) in thirty DSFGs at .
+The AS2VLA survey used JVLA to observe the CO(1–0) emission (rest-frame frequency: = 115.2712 GHz) in thirty DSFGs at .
 The surveys comprises of the JVLA programs ID:21A-254 (PI: Hodge) and ID:17A-251 (PI: Walter).
@@
 2007) versions 6.4.1 and 6.4.3 and the standard JVLA pipeline.
-The observations of the first 17 galaxies (eleven secure detections) were presented in Frias Castillo et al.
-(2023).
-In this work (Paper II), we focus on the 19 sources with secure CO(1–0) detections333For a “secure” detection, we require peak S/N3 in naturally-weighted cleaned maps of CO(1–0) emission.
-from the completed survey (Paper III, in prep).
+The observations of the first 17 galaxies (11 robust detections) were presented in Frias Castillo et al.
+(2023, Paper I).
+The remaining sources from the completed survey will be presented in a forthcoming paper (Paper III); out of these, an additional eight sources were found to have secure detections of the CO(1-0) emission line.
+For this work (Paper II), we selected all the sources of the entire sample having secure CO(1-0) detections (peak S/N3 in naturally-weighted cleaned maps of CO(1-0) emission), amounting to a total of 19 sources.
 2.3 Visibility-plane analysis and stacking
@@
 1.
+We achieve an rms sensitivity of 0.014 Jy km s-1 beam-1 with a median beam FWHM 4.2”2.7”, corresponding to K km s-1 pc2 kpc-2 at , the median redshift of our sample.
 Image-plane deconvolution with the imfit task in Casa does not yield a reliable source size.
@@
 The fields were centred at the ALMA dust continuum peak (accurate to 0.1).
-We confirmed that the phase-tracking centre is well-aligned with the CO(1–0) emission; we phase-shifted three sources with noticeable offsets (AEG2, CDFN1, CDFN2).
+The phase-tracking centre is generally well-aligned with the CO(1–0) emission; we phase-shifted three sources with noticeable offsets (AEG2, CDFN1, CDFN2).
 For two sources (AS2COS14.1, AS2COS31.1), we found a bright continuum source within the VLA primary beam.
@@
 We fit each binned visibility function with: (1) point-source (constant signal) and (2) an exponential profile.
-We prefer the exponential profile to the more widely-used Gaussian profile as it provides a more physical description of gas surface brightness distribution in star-forming galaxies 444For the literature source sizes derived using Gaussian profiles, we convert the half-light radius from the Gaussian to exponential profile.
+We prefer the exponential profile to the more widely-used Gaussian profile as it provides a more physical description of gas surface brightness distribution in star-forming galaxies 333For the literature source sizes derived using Gaussian profiles, we convert the half-light radius from the Gaussian to exponential profile.
 (e.g., Leroy et al.
@@
 2024.
-We utilise the Monte-Carlo Markov chain modelling using the emcee package (Foreman-Mackey et al., 2013).
+We utilise Monte-Carlo Markov chain modelling using the emcee package (Foreman-Mackey et al., 2013).
 We assumed wide, uniform priors in total flux and half-light radius.
 For each galaxy, we calculate the Akaike and Bayesian information criteria (AIC, BIC) for each model.
-The results for individual galaxies are listed in Table 2.
+Table 1 lists the source sizes inferred from the MCMC modelling.
+For galaxies for which the extended source is not preferred by both AIC and BIC, we give upper limit on the source sized, based on the 95th percentile from the posterior of the exponential model.
+Source ID | CO(1–0) | [C i](1–0) | |||
+[arcsec] | [kpc] | [arcsec] | [kpc] | ||
+AS2COS1.1 | 4.625 | 2.62 | 17 | 1.36 | 9 |
+AS2COS8.1 | 3.581 | 1.05 | 8 | 0.75 | 6 |
+AS2COS13.1 | 2.608 | 0.41 | 2.9 | — | — |
+AS2COS14.1 | 2.921 | 0.74 | 5.8 | 0.98 | 7.7 |
+AS2COS23.1 | 4.341 | 1.89 | 13 | 1.110.44 | 7.63.0 |
+AS2COS28.1 | 3.097 | 0.66 | 5.1 | — | — |
+AS2COS31.1 | 3.643 | 5.26 | 39 | 0.44 | 3.4 |
+AS2COS54.1 | 3.174 | 1.60 | 12 | — | — |
+AS2COS65.1 | 2.414 | 0.68 | 5.6 | 0.380.11 | 3.10.9 |
+AS2COS66.1a | 3.247 | 2.09 | 17 | — | — |
+AS2COS139.1 | 2.219 | 1.08 | 8 | — | — |
+AS2UDS10.0 | 3.169 | 3.65 | 28 | — | — |
+AS2UDS11.0 | 4.073 | 2.37 | 38 | — | — |
+AS2UDS12.0 | 2.520 | 0.80 | 6.6 | 0.400.19 | 3.31.6 |
+AS2UDS126.0 | 2.436 | 0.97 | 8.0 | 0.640.28 | 5.32.3 |
+AEG2 | 3.668 | 1.43 | 12 | — | — |
+CDFN1 | 3.149 | 0.80 | 6 | — | — |
+CDFN2 | 4.422 | 1.35 | 9 | — | — |
+CDFN8 | 4.144 | 0.89 | 6 | — | — |
+AS2COS2.1 | 4.595 | — | — | 2.30 | 15 |
+AS2UDS.26 | 3.296 | — | — | 0.67 | 5.1 |
+a The [C i] line in AS2COS66.1 falls partially outside the ALMA tuning | |||||
+and is not included in this analysis. |
 3 Analysis and results
 Fits to individual galaxies show that the point-source model (our null hypothesis) is preferred for 15 out of 19 sources.
-For four sources, exponential profiles with are preferred (see Table 2): these are AS2COS13.1, AS2COS14.1, AS2COS28.1555AS2COS28.1 has a companion offset by 20 kpc, which might contribute to the large apparent source size (Chen et al., 2022a)., and AS2UDS126.0.
+For four sources, exponential profiles with are preferred (see Table 1): these are AS2COS13.1, AS2COS14.1, AS2COS28.1444AS2COS28.1 has a companion offset by 20 kpc, which might contribute to the large apparent source size (Chen et al., 2022a)., and AS2UDS126.0.
 These four sources do not stand out from the rest of the sample in redshift, SFR or .
@@
 Are the extended CO reservoirs associated with molecular outflows or inflows?
-Molecular outflows – indicated by high-velocity wings of the line profile - have been invoked to explain extended [C ii] reservoirs around high-redshift galaxies from the ALPINE sample (Ginolfi et al., 2020; Pizzati et al., 2020, 2023); while ALPINE galaxies have SFRs an order-of-magnitude lower than our sample (10–100 M⊙/yr), a similar mechanism might be at play at even more highly star-forming DSFGs.
+Gas outflows – indicated by high-velocity wings of the line profile – have been invoked to explain extended [C ii] reservoirs around high-redshift galaxies from the ALPINE sample (Ginolfi et al., 2020; Pizzati et al., 2020, 2023).
+Although ALPINE galaxies have SFRs an order-of-magnitude lower than our sample (10–100 M⊙ yr-1), the same mechanism might be at play at even more highly star-forming DSFGs.
 To search for potential high-velocity components, we stack the spectra of all 19 CO(1–0)-detected galaxies.
@@
 As the half-light radius of the stacked emission is smaller than the VLA beam, we extract spectra from an aperture equal to the beam size.
-The stacked spectrum is well-fit by a single Gaussian profile with a line FWHM of 64070 km s-1 (Fig.
+The stacked spectrum is well-fit by a single Gaussian profile with a line FWHM of 59070 km s-1 (Fig.
 2).
@@
 This residual continuum emission is likely due to curvature of the spectrum: at the rest-frame 115 GHz, the combination of free-free and dust thermal emission can lead to a significant spectral curvature.
-The lack of obvious outflow features indicates that the bulk of CO(1–0) gas is close to the systemic velocity.
+The lack of obvious outflow features indicates that the bulk of CO(1–0) is close to the systemic velocity.
 3.2 ALMA [CI](1–0) data
-We supplement our CO(1–0) data with [C i](1–0) ALMA observations of seven galaxies from our sample which were recently detected in [C i](1–0) line by Frias Castillo et al.
-(2024).
+We supplement our CO(1–0) data with [C i](1–0) ALMA observations of ten galaxies from our parent sample detected in [C i](1–0) line by Chen et al.
+(2022a, AS2COS31.1) and Frias Castillo et al.
+(2024, nine galaxies555The tenth galaxy detected in () – AS2COS66.1 – has only part of the [C i](1–0) line covered by the ALMA tuning; consequently, we do not include it in our analysis).
 The [C i](1–0) line is considered an alternative tracer of molecular gas, including the “CO-dark” phase (e.g., at low metallicities).
@@
 The two sources that are extended in both CO(1–0) and [C i](1–0) – AS2COS14.1 and AS2UDS126.0 – have consistent sizes in both tracers.
-Stacking all the data, we find that the exponential profile with =0.280.10 (2.20.8 kpc) is preferred over the point-source, albeit at lower significance than for CO(1–0) (AIC2.5).
-The inferred mean [C i](1–0) size is 4025% smaller than the CO(1–0) one, although they are consistent within 1.7.
-We therefore put an upper limit on the [C i](1–0) size by taking the 95th percentile of the posterior ( 0.42 (3.3 kpc).
-We thus can not rule out the that [C i](1–0) and CO(1–0) have similar scale lengths.
+Stacking all the data, we find that the exponential profile with =0.200.11 (1.60.8 kpc) is preferred over the point-source one, albeit at low significance (AIC2).
+The inferred mean [C i](1–0) size is 6025% smaller than the CO(1–0) one, as a 2.3 significance.
+Given the very weak preference for the exponential profile over the point-source one, we conservatively put an upper limit on the [C i](1–0) size by taking the 95th percentile of the posterior ( 0.33 (2.6 kpc).
+The [CI](1-0) therefore appears to be marginally less extended than CO(1-0).
 3.3 PDR modelling
@@
 Specifically, we estimate the gas density and far-UV irradiation using the PDRToolbox suite of uniform-density PDR models (Kaufman et al., 2006; Pound & Wolfire, 2008).
-As we will show in Section 4, the CO(1–0)/[C i](1–0) sizes inferred from the uv-plane stacking are consistent with the extended [C ii] and far-IR continuum emission seen in previous studies of DSFGs.
+As we will show in Section 4, the CO(1–0)/[C i](1–0) sizes inferred from the uv-plane stack are consistent with the extended component of the [C ii] and far-IR continuum emission seen in previous studies of DSFGs.
 For the far-IR continuum, we assume that the extended component accounts for 13% of the total luminosity (Gullberg et al.
@@
 To assess the sensitivity of our conclusions to the various assumptions made in the PDR modelling we have varied these as follows:
-What if CO(1–0) is suppressed by CMB? Assuming a 50% suppression (an extreme scenario at , e.g., da Cunha et al. 2013), the inferred far-UV illumination increases by 0.5 dex (blue dotted line).
 What if the bulk of [C ii] is arising from ionised gas rather than from the PDRs?
-Lowering the PDR contribution to the [C ii] flux down to 20% causes the inferred far-UV illumination to increase by almost 1 dex (red dotted line).
+Lowering the PDR contribution to the [C ii] flux down to 20% causes the inferred far-UV illumination to increase by almost 1 dex (red dotted line in Fig.
+4).
 What if [C i](1–0) is more compact than CO(1–0) (i.e., the ratio of atomic/molecular carbon increases at large radii)?
 We assume the [C i](1–0) scale length indicated by the tentative results in Section 3.2 and re-calculate the [C i](1–0) surface brightness at the CO(1–0) half-light radius accordingly.
-The inferred density increases to cm-3 (green dotted line).
+The inferred density increases to cm-3 (green dotted line in Fig.
+4).
 We derive the filling factor – i.e., the number of emitting regions per line-of-sight – by comparing the observed CO(1–0) surface brightness with the PDRToolbox models.
 Specifically, for each realisation of , , we calculate the corresponding filling factor at the CO(1–0) half-light radius as:
-where is the CO(1–0) flux per solid angle predicted by PDRToolbox and is the luminosity distance. Marginalising over and then yields %.
+where is the CO(1–0) flux per solid angle predicted by PDRToolbox and is the luminosity distance. Marginalising over and then yields %. We discuss the context of this result in Section 4.4.
 4 Discussion
+In this Section, we compare the results of our stacking procedure and PDR modelling to other studies of high-redshift DSFGs (4.1), dust and stellar emission (4.2), and hydrodynamical simulations (4.3).
+Finally, we address the nature of extended CO(1–0) reservoirs (4.4).
 4.1 Sizes of extended molecular gas reservoirs in high-redshift DSFGs
@@
 Specifically, we show the literature CO(1–0) (Ivison et al., 2011; Thomson et al., 2012; Dannerbauer et al., 2017; Frias Castillo et al., 2022; Stanley et al., 2023) and [C ii] measurements (Gullberg et al., 2018; Rybak et al., 2019; Mitsuhashi et al., 2021).
-We homogenise the data by adopting the empirical CO excitation ladder from Birkin et al.
-(2021) and ; for the [C ii] measurements, we assume =30 from Zanella et al.
+We derive with (e.g., Calistro Rivera et al.
+2018).
+For the [C ii]-based masses, we use , assuming =30 (Zanella et al., 2018).
+Our inferred CO(1–0) size is comparable to CO(1–0) sizes reported in the literature for other highly star-forming galaxy populations, and the [C ii] reservoirs seen around high-redshift DSFGs.
+Our inferred sizes are also consistent with the CO(1–0) sizes reported by Stanley et al.
+(2023) for the (potentially) non-lensed sources from the Z-GAL sample (median 4.4 kpc).
+In contrast, our CO(1–0) size is larger than the typical extent of CO(3–2) emission in DSFGs—2.6 kpc—recently reported by Amvrosiadis et al.
 (2025).
-Our inferred CO(1–0) size is comparable to CO(1–0) sizes reported in the literature for other highly star-forming galaxy populations, and the [C ii] reservoirs seen around high-redshift DSFGs.
 The true extent of the molecular gas reservoirs in DSFGs might be even larger than those inferred from our CO(1–0) stacking.
@@
 However, the CMB effect will be more limited if the molecular gas is mostly in the form of higher-density ISM in star-forming regions or satellite galaxies.
+The CO(1–0) extent derived from our stacking is comparable to the sizes of [C ii] emission in high-redshift galaxies with similar SFR and molecular gas mass (Fig.
+5, e.g., Gullberg et al.
+2018; Rybak et al.
+2019; Mitsuhashi et al.
+2021), as well as those reported around star-forming galaxies from the CRISTAL sample ( kpc, Ikeda et al.
+2025).
+The extended [C ii] reservoirs around high-redshift galaxies have been hypothesised to arise primarily from atomic or ionised gas (e.g., Ikeda et al.
+2025), perhaps deposited by outflows (e.g., Pizzati et al.
+2020, 2023.
+However, the comparable extent of the CO(1–0) (which is associated only with molecular gas) indicates that molecular gas might be the dominant component of the extended gas halos.
 4.2 Molecular gas versus dust and stellar emission
@@
 (2019) found a half-light radius of 1.10.3 kpc, assuming an exponential profile666After applying a correction to account for a faint extended component (Smail et al., 2021)..
-Somewhat larger sizes (1.80.1 kpc) were inferred for the ALESS sample of DSFGs (Hodge et al., 2016, 2024), which span a comparable redshift range (median ) and 850-m flux ( mJy) to our galaxies.
-In either case, the 870-m continuum is a factor of a more compact than the CO(1–0) reservoir.
+Somewhat larger sizes (1.80.1 kpc) were inferred for the ALESS sample of DSFGs (Hodge et al., 2016, 2025), which span a comparable redshift range (median ) and 850-m flux ( mJy) to our galaxies.
+In either case, the 870-m continuum is a factor of more compact than the CO(1–0) reservoir.
 In total, 50–80% of the CO(1–0) emission arises outside the region which contains 90% of the dust continuum emission (adopting the Gullberg et al.
@@
 The presence of an extended molecular gas dovetails with the extended low-surface brightness dust continuum reported by Gullberg et al.
-(2019).
-After stacking ALMA imaging of 153 DSFGs, Gullberg et al.
-(2019) reported that 13% of the 870-m continuum arises for an extended component with an assumed radius of 4 kpc, similar to our inferred CO(1–0) size.
+(2019), who after stacking ALMA imaging of 153 DSFGs, reported that 13% of the 870-m continuum arises from an extended component with an assumed radius of 4 kpc, which is similar to our measured CO(1–0) size.
 This extended continuum emission lends further support to the presence of significant metal-enriched gas and reservoirs around high-redshift DSFGs.
-As the bulk of molecular gas is far outside the central star-forming region, can the high SFRs of DSFGs be sustained by smooth gas accretion from this reservoirs?
-We can estimate the required inward streaming velocity by dividing the half-light radius by the median gas depletion time ( Myr, Frias Castillo et al.
+As the bulk of the molecular gas is far outside the central star-forming region, can the high SFRs of DSFGs be sustained by smooth gas accretion from this reservoirs?
+We can estimate the required inward streaming velocity by dividing the half-light radius by one half of the median gas depletion time ( Myr, Frias Castillo et al.
 2023).
-We obtain km s-1, comparable to the line-of-sight velocity dispersions seen in high-resolution kinematic studies of DSFGs (10 - 100 km s-1, e.g., Hodge et al.
+We obtain km s-1, consistent with to the line-of-sight velocity dispersions seen in high-resolution kinematic studies of DSFGs (10 - 100 km s-1, e.g., Hodge et al.
 2012; Rybak et al.
@@
 2021, 2023; Birkin et al.
-2024.
-The central star-formation can thus be supported by smooth streaming of gas within the galaxy without major mergers (see, e.g., Narayanan et al.
+2024; Amvrosiadis et al.
+2025.
+The central star formation can thus be supported by smooth streaming of gas within the galaxy without major mergers (see, e.g., Narayanan et al.
 2015).
@@
 (2025).
-The CO(1–0) emission in our sample is a factor of 1.40.2 and 1.30.2 more extended than 4.4-m sizes from Gillman et al.
+The CO(1–0) emission in our sample is a factor of 1.40.2 and 1.30.2 more extended than the 4.4-m sizes from Gillman et al.
 (2024) and Hodge et al.
@@
 CO(1–0) | 3.80.5 | This work |
-[C i](1–0) | 3.3 | This work |
+[C i](1–0) | 2.6 | This work |
 870 m | 1.10.3 | Gullberg et al. (2019) |
@@
 Second, if the dust temperature decreases towards larger radii (e.g., Calistro Rivera et al.
-2018) star-formation will be more centrally concentrated than the observed-frame 870-m continuum.
-Finally, stellar component is likely more compact than the near-infrared continuum; high obscuration in the central region will “inflate” the apparent near-infrared sizes compared to the actual stellar component (Sorba & Sawicki, 2018; Popping et al., 2022; Smail et al., 2023; Gillman et al., 2024).
+2018) star formation will be more centrally concentrated than the observed-frame 870-m continuum.
+Finally, the stellar component is likely more compact than the near-infrared continuum; high obscuration in the central region will “inflate” the apparent near-infrared sizes compared to the actual stellar component (Sorba & Sawicki, 2018; Popping et al., 2022; Smail et al., 2023; Gillman et al., 2024).
 How do the relative sizes of molecular gas, stars and dust continuum in high-redshift DSFGs compare to other galaxy populations?
@@
 SIMBA successfully reproduces key parameters of high-redshift DSFGs (Lovell et al., 2021).
-We focus the snapshot from the default 100-cMpc box.
+We focus on the snapshot from the default 100-cMpc box.
 We select galaxies with M⊙, i.e., comparable in mass to our sample.
@@
 2019).
-4.4 The nature and thermodynamics of the extended CO(1–0) reservoirs
+4.4 The nature of the extended CO(1–0) reservoirs
 What is the nature of the extended CO(1–0) emission around DSFGs?
@@
 2024; Hodge et al.
-2024).
-In either scenario, the molecular gas on the galaxy outskirts is significantly clumpy rather than distributed in a smooth, diffuse “halo”.
+2025).
+In either scenario, the molecular gas on the galaxy outskirts would be significantly clumpy rather than distributed in a smooth, diffuse “halo”.
 5 Conclusions
 We have presented results of stacking CO(1–0) observations of a sample of 19 high-redshift DSFGs.
-Comprising 80 hours in the most compact JVLA array configuration, this is the deepest such study to date.
+Comprising 80 hours in the most compact JVLA array configuration, this is the deepest study of extended cold-gas reservoirs around early galaxies to date.
 Our main findings are:
-The stacked -plane data show an extended exponential profile with a half-light radius of 3.80.5 kpc.
+The stacked -plane data of the CO(1–0) emission show an extended exponential profile with a half-light radius of 3.80.5 kpc.
 The inferred size is comparable to extended CO(1–0) and [C ii] reservoirs previously reported around high-redshift DSFGs.
 Four galaxies have clearly extended CO(1–0) emission, with half-light radii between 3.4–7.2 kpc.
-Applying the same procedure to ALMA [C i](1–0) observations, we find that [C i](1–0) appears marginally more compact (2.20.8 kpc) than CO(1–0), although both sizes are consistent within 2.
+Applying the same procedure to ALMA [C i](1–0) observations, we find that [C i](1–0) appears more compact (2.6 kpc) than CO(1–0), although both sizes are consistent within 2.
+Five galaxies show extended [C i](1–0) emission, with half-light radii of 3.1–7.7 kpc.
 Using PDR modelling to infer physical conditions in the extended CO(1–0) reservoirs, we find a density of cm-3 and a far-UV field of , and a filling factor of 1%.
 These results indicate that the extended CO (and [C i] and [C ii]) emission arises from sparsely distributed star-forming clouds.
-The inferred sizes of CO(1–0) reservoir are significantly larger than the 870-m dust continuum, indicating that the bulk of molecular gas (50–80%) resides outside the FIR-bright star-forming region.
-CO(1–0) is somewhat larger than the 4.4-m stellar emission.
-Ultimately, understanding the nature of extended molecular gas reservoirs around high-redshift galaxies will require a new class of large, single-dish telescopes, such as the proposed 50-metre Atacama Large Aperture Submillimeter Telescope (AtLAST) (Mroczkowski et al., 2024; Lee et al., 2024).
+The inferred sizes of CO(1–0) cold-gas reservoir are significantly larger than the bulk of the 870-m dust continuum, but similar to the fainter, extended dust component that has been reported in these sources.
+The CO(1–0) is somewhat larger than the 4.4-m stellar emission.
+The bulk of the molecular gas (50–80%) resides outside the FIR-bright star-forming region and does not directly contribute to the observed star formation.
+Our findings reveal that the bulk of molecular gas in high-redshift DSFGs resides in extended, 10-kpc scale reservoirs; only a small fraction of this gas directly contributes to the observed star formation.
 Acknowledgements.
@@
 Pizzati and A.
-Schimek for sharing insights from theoretical models and simulations.References
+Schimek for sharing insights from theoretical models and simulations.
+Finally, we thank the referee – P.
+Cox – for his detailed comments that helped us improve this manuscript.References
+- Amvrosiadis et al. (2025) Amvrosiadis, A., Wardlow, J. L., Birkin, J. E., et al. 2025, MNRAS, 536, 3757
+- Battisti et al. (2019) Battisti, A. J., da Cunha, E., Grasha, K., et al. 2019, ApJ, 882, 61
 - Bellocchi et al. (2022) Bellocchi, E., Pereira-Santaella, M., Colina, L., et al. 2022, A&A, 664, A60
@@
 - da Cunha et al. (2013) da Cunha, E., Groves, B., Walter, F., et al. 2013, ApJ, 766, 13
+- da Cunha et al. (2015) da Cunha, E., Walter, F., Smail, I. R., et al. 2015, ApJ, 806, 110
 - Dannerbauer et al. (2017) Dannerbauer, H., Lehnert, M. D., Emonts, B., et al. 2017, A&A, 608, A48
@@
 - Frayer et al. (2011) Frayer, D. T., Harris, A. I., Baker, A. J., et al. 2011, ApJ, 726, L22
+- Frayer et al. (2018) Frayer, D. T., Maddalena, R. J., Ivison, R. J., et al. 2018, ApJ, 860, 87
 - Frias Castillo et al. (2023) Frias Castillo, M., Hodge, J., Rybak, M., et al. 2023, ApJ, 945, 128
@@
 - Hodge & da Cunha (2020) Hodge, J. A. & da Cunha, E. 2020, Royal Society Open Science, 7, 200556
-- Hodge et al. (2024) Hodge, J. A., da Cunha, E., Kendrew, S., et al. 2024, arXiv e-prints, arXiv:2407.15846
+- Hodge et al. (2025) Hodge, J. A., da Cunha, E., Kendrew, S., et al. 2025, ApJ, 978, 165
 - Hodge et al. (2016) Hodge, J. A., Swinbank, A. M., Simpson, J. M., et al. 2016, ApJ, 833, 103
+- Ikeda et al. (2025) Ikeda, R., Tadaki, K.-i., Mitsuhashi, I., et al. 2025, A&A, 693, A237
 - Ivison et al. (2011) Ivison, R. J., Papadopoulos, P. P., Smail, I., et al. 2011, MNRAS, 412, 1913
@@
 - Kaufman et al. (2006) Kaufman, M. J., Wolfire, M. G., & Hollenbach, D. J. 2006, ApJ, 644, 283
-- Lee et al. (2024) Lee, M. M., Schimek, A., Cicone, C., et al. 2024, arXiv e-prints, arXiv:2403.00924
 - Leroy et al. (2009) Leroy, A. K., Walter, F., Bigiel, F., et al. 2009, AJ, 137, 4670
@@
 - Mitsuhashi et al. (2021) Mitsuhashi, I., Matsuda, Y., Smail, I., et al. 2021, ApJ, 907, 122
-- Mroczkowski et al. (2024) Mroczkowski, T., Gallardo, P. A., Timpe, M., et al. 2024, arXiv e-prints, arXiv:2402.18645
 - Narayanan et al. (2015) Narayanan, D., Turk, M., Feldmann, R., et al. 2015, Nature, 525, 496
@@
 - Zhang et al. (2016) Zhang, Z.-Y., Papadopoulos, P. P., Ivison, R. J., et al. 2016, Royal Society Open Science, 3, 160025
-Appendix A Inferred source sizes
-Table 2 lists the source sizes inferred from the MCMC modelling, using 5-k radial bin size.
-For galaxies for which the extended source is not preferred by both AIC and BIC, we give upper limit on the source sized, based on the 95th percentile from the posterior of the exponential model.
-Source ID | CO(1–0) | [C i](1–0) | |||
-[arcsec] | [kpc] | [arcsec] | [kpc] | ||
-AS2COS1.1 | 4.625 | 2.62 | 17 | 1.36 | 9 |
-AS2COS8.1 | 3.581 | 1.05 | 8 | 0.75 | 6 |
-AS2COS13.1 | 2.608 | 0.41 | 2.9 | — | — |
-AS2COS14.1 | 2.921 | 0.74 | 5.8 | 0.98 | 7.7 |
-AS2COS23.1 | 4.341 | 1.89 | 13 | 1.110.44 | 7.63.0 |
-AS2COS28.1 | 3.097 | 0.66 | 5.1 | — | — |
-AS2COS31.1 | 3.643 | 5.26 | 39 | — | — |
-AS2COS54.1 | 3.174 | 1.60 | 12 | — | — |
-AS2COS65.1 | 2.414 | 0.68 | 5.6 | 0.380.11 | 3.10.9 |
-AS2COS66.1 | 3.247 | 2.09 | 17 | — | — |
-AS2COS139.1 | 2.219 | 1.08 | 8 | — | — |
-AS2UDS10.0 | 3.169 | 3.65 | 28 | — | — |
-AS2UDS11.0 | 4.073 | 2.37 | 38 | — | — |
-AS2UDS12.0 | 2.520 | 0.80 | 6.6 | 0.400.19 | 3.31.6 |
-AS2UDS126.0 | 2.436 | 0.97 | 8.0 | 0.640.28 | 5.32.3 |
-AEG2 | 3.668 | 1.43 | 12 | — | — |
-CDFN1 | 3.149 | 0.80 | 6 | — | — |
-CDFN2 | 4.422 | 1.35 | 9 | — | — |